Calculate the entropy, for the freezing of 20.0g liquid metal mercury at -38.9˚C.(molar enthalpy of fusion of mercury is 2.33kJ/mol).


1
Expert's answer
2021-06-14T13:34:49-0400

Entropy = heat released/ temperature (K)

temperature = -38.9⁰C = 234.1 K

heat released = no. of moles × "\Delta H_{fus}"

no of moles = 20 g/200gmol-1 = 0.1 mol

heat released = 0.1 × 2.33 = 0.233KJ

Entropy = 0.233 / 234.1 = 0.0009953 KJ/K